OpenStudy (anonymous):

Solve 2 log2 2 + 2 log2 6 – log2 3x = 3

OpenStudy (anonymous):

the log(4)(36)-log36=3

OpenStudy (anonymous):

hmm ,from where you got (36) ???

OpenStudy (anonymous):

i meant to put log(4)(36)-log3x=3 sorry

OpenStudy (anonymous):

right now i really just don't know where to go from log144-log3x=3

OpenStudy (anonymous):

Wow I would have never gotten that thank you! but thats not one of the answer choices

OpenStudy (anonymous):

ok im confused

OpenStudy (anonymous):

thay are x= 2,x=6,x=16,x=18

OpenStudy (anonymous):

Lets begin step by step First we will get the power inside the log : \[\Large Log_{2}2^2+Log_{2}6^2-Log_{2}(3x)=3\] Then according the the property that i showed to you in the last Post We will combine them: \[\Large Log_{2}\frac{2^2*6^2}{(3x)}=3\] Lets change it to exponential : \[\Large \frac{2^3}{1}=\frac{2^2*6^2}{(3x)}\] Now lets change it to an equation: \[\Large \color{red}{24x=144}\] Divide by 24 \[\Huge \color{blue}{x=6}\]

OpenStudy (anonymous):

Got it ?

OpenStudy (anonymous):

ooh ok i got it thanks